What is the length of a ramp that has a 25.1 degree incline and rises 11.5 m?
1
Expert's answer
2019-05-09T11:58:56-0400

Since the height is 11.5 m, calculate the length of the bottom of the ramp:


"\frac{\text{height}}{\text{length}}=\text{tan}\alpha,"

"\text{length}=\frac{\text{height}}{\text{tan}\alpha}=\frac{11.5}{\text{tan}25.1^\circ}=24.5\text{ m}."